|
Документ взят из кэша поисковой машины. Адрес
оригинального документа
: http://www.stsci.edu/smov/notes/0223smov.html
Дата изменения: Tue Mar 11 23:58:55 1997 Дата индексирования: Thu Feb 27 23:53:55 2014 Кодировка: Поисковые слова: http www.badastronomy.com phpbb index.php |
STATUS REPORTS
1. Observatory Status
MOSES status reports:
-PCS - nominal - Proposal 7028 executed successfully,
returning HST to fine-lock guiding (FGS 2&3)
`
-EPS - nominal - SADE timing will continue to be monitored.
-DMS - nominal - SSR is currently continuously recording
engineering data (SMS053).
-Thermal - nominal.
-FSW - STIS 4.5 patch installation was carried out.
(Hand-out provided by SI/MOSES.) STIS was
returned to HOLD mode to await SMS reconfiguration.
Post-installation EEPROM memory dump was verified by
BRZ updated to match on-orbit SI configuration.
2. WFPC2 Status - nominal - successfully transitioned from SAFE
to HOLD. CCDs at +22C. AFMs at -43C.
3. NICMOS Status - During execution of proposal 7033, two status
buffer messages were issued due to an improperly commanded
NICMOS memory monitor word during the memory self test. The
second was due to a memory dump time-out because of improper
SDF conditioning by the SMS commands for dumping. (Details
in hand-outs provided by both STScI and MOSES.) The
NICMOS EEPROM and CS buffer memory were dumped successfully
in real-time. Now nominal.
4. STIS Status - During proposal 7058 (STIS Memory Load & Dump),
the instrument suspended, reporting two status buffe messages.
Both were related to the attempt to invoke an OPERATE-mode
macro while in BOOT mode. The SI was recovered by PSTOL
commands. (Details in a MOSES-provided hand-out.)
Now nominal, although the IDT reports that 28-volt noise
problem has reappeared without any reconfiguration of FOC/PDU
(See notes from 22 Feb meeting.) Believed to be understood
as some sort of amplification effect.
5. FGS status - nominal - no change from last report
6. FOC/COSTAR status - COSTAR DOB was successfully deployed, but not
before a safing event caused by a mis-match between actual
start position and the position expected by new NSSC1 FSW.
(Details in STScI-provided handout.)
PLANNING REPORTS
1. Status of Upcoming SMSs
- SMS053 (1st SMOV Science SMS) - Uplinked and
executing.
The goal is to run all of the SMSs through SMS069 by
Sunday for input to commanding review in order to
get a jump on TDRS requests in forecast period.
- SMS055 - Signed off. Ready for uplink.
- SMS058 - Has beem built and finals have been delivered
for review.
- SMS058H - (contingency SMS058 in case WFPC2
contamination is seen in SMS055)
- Has beem built and finals have been delivered
for review.
- SMS062 - built and under commanding review.
TDRSS requests being negotiated Monday morning during
Active Period.
- SMS065 - built and under commanding review. TDRSS requests
being negotiated Monday morning.
- SMS069 - built and under commanding review. TDRSS requests
being negotiated Monday morning prior to forecast
period.
- SMS072 - currently being built, expected tonight for initial
SARs.
2. Replans - none
OPERATIONAL ISSUES - As a result of the STIS and NICMOS operational problems
reported above, John Campbell requests a list
from the STScI of all SMOV proposals using special commanding and
requests a furhter review, on the part of the teams, of the upcoming
activities in the current and next SMSs.